OpenStudy (booksforlife):

A jet ski depreciates at 11% of its original value each year. If the jet ski was $8,000 at its time of purchase, what is the value of the jet ski after 5 years?

OpenStudy (booksforlife):

@SaiyanSparta could you help me with this one?

OpenStudy (saiyansparta):

Since the ski depreciates at 11% of its original value each year, after one year, it will be at 100-11 = 89%, or 0.89 of its value from the previous year. Therefore, in 5 years, it will be \[(0.89)^5\] of its original value. Therefore, the value is \[(0.89)^5\cdot \$ 8,000 \approx \$ 4,467.25 \]

OpenStudy (booksforlife):

Thank you again @SaiyanSparta! This really helped a lot! :)

OpenStudy (saiyansparta):

Anytime! :)

OpenStudy (booksforlife):

Could you check my answer on this one? 1/36=6^(x-3) 1/6^2 = 6^(1)(x-3) 6^-2=6^(x-3) -2=x-3 -2+3=x-3+3 1=x

OpenStudy (booksforlife):

@SaiyanSparta sorry if I'm asking a lot. I've been stuck on these for 3 days.

OpenStudy (saiyansparta):

Yup that seems about right! :)